OpenStudy (anonymous):

mc003-1.jpg If AD = 3x + 3 and CD = x + 9, what is the value of x?

OpenStudy (anonymous):

send the pic

OpenStudy (anonymous):

post the figure please....

OpenStudy (anonymous):

use AD=CD 3x+3=x+9 i hop e u can solve now

OpenStudy (anonymous):

AD is marked as congruent to CD AD = CD 3x+3=x+9 2x+3 = 9 2x=6 x=?